OpenStudy (anonymous):

For the curve y=x^3/12+x^-1 I must find the length from x=1 to x=3.

OpenStudy (anonymous):

To be more specific, I don't need any sort of answer, just a bit of help getting past one part of it so I can finish it on my own. So, the arclength is equal to \[\int\limits_{a}^{b}\sqrt{1+(\frac{ dy }{ dx })^2 dx}\] I end up getting dy/dx=1/4x^2+1/x^2, which when squared is equal to x^4/16+1/2+1/x^4. Putting that back in it gives us \[\int\limits_{a}^{b}\sqrt{\frac{ x^4 }{ 16 }+ \frac{ 3 }{ 2 }+\frac{ 1 }{ x^4 }} dx\] I don't know how to manipulate it into anything easier (I have to integrate it by hand.)

OpenStudy (irishboy123):

\[ y=\frac{x^3}{12}+\frac{1}{x}\] \[y' = \frac{x^2}{4} - \frac{1}{x^2}\]

OpenStudy (irishboy123):

ie \[y' = \frac{x^2}{4} \color{red}- \frac{1}{x^2}\]

OpenStudy (anonymous):

I had that originally but even then it was still a really ugly trinomial.

OpenStudy (irishboy123):

\[y'^2 = \frac{x^4}{16} - \frac{1}{2} + \frac{1}{x^4}\] \[y'^2 \color{red}{+ 1 }= \frac{x^4}{16} + \frac{1}{2} + \frac{1}{x^4}\] \[= (\frac{x^2}{4} + \frac{1}{x^2})^2\]

OpenStudy (anonymous):

Yeah, so that's what I had down on my paper but didn't realize it would make a perfect square. Thank you sir.

OpenStudy (irishboy123):

mp!